OpenStudy (anonymous):

y varies directly as the square of x. When x = 3, y = 54. Find y when x = 8. y = 96 y = 144 y = 256 y = 384

OpenStudy (jhannybean):

If y varies directly we can follow the format of direct variation as \[\large y=kx^2\]\[\large x=3 \ ,\ \ y=54\]first we solve for k.\[\large 54=(3)^2k\]\[\large 54 = 9k\]\[\large k=6\] Now we're asked to find y when \(x=8\) \[\large y=(8)(6)\]\[\large y =?\]

OpenStudy (anonymous):

12

OpenStudy (anonymous):

48

OpenStudy (anonymous):

srry

OpenStudy (jhannybean):

Sorry I, too, made a mistake.

OpenStudy (anonymous):

288

OpenStudy (anonymous):

y varies directly as the square of x. When x = 3, y = 54. Find y when x = 8. y = 96 y = 144 y = 256 y = 384

OpenStudy (jhannybean):

One more time! lol. \[\large y=kx^2\]\[\large k= 6 \ ,\ x=8\]\[\large y=(6)(8)^2\]\[\large y= 6 \times 64\]\[\large y=384\] I had mixed up my k and my x. :(